OpenStudy (anonymous):

Write the equation x^2 + 7y^2 - 12x - 13 in standard form. State which type of conic section it is. Please Help Medals will be given.

OpenStudy (jmark):

Given x^2 + 7y^2 - 12x - 13 Both the x^2 and y^2 are present so it is ellipse or hyperbola The sign of the coefficients are the same so it is an ellipse. Complete the square on the x and on the y terms. x^2-12x +36 + 7(y^2) = 13+36 Factor: (x-6)^2 + 7(y-0)^2 = 49 Standardize (x-6)^2/7^2+(y-0)^2/(sqrt(7))^2 = 1 Center: (6,0) Semi major axis: 7 (horizontal) ICSE sample papers - http://icse.edurite.com/icse-sample-papers.html